# SpatialRGPT: Grounded Spatial Reasoning in Vision-Language Models

### v2

@@ Abstract
Vision Language Models (VLMs) have demonstrated remarkable performance in 2D vision and language tasks. However, their ability to reason about spatial arrangements remains limited. In this work, we introduce Spatial Region GPT (SpatialRGPT) to enhance VLMs’ spatial perception and reasoning capabilities. SpatialRGPT advances VLMs’ spatial understanding through two key innovations: (i) a data curation pipeline that enables effective learning of regional representation from 3D scene graphs, and (ii) a flexible “plugin” module for integrating depth information into the visual encoder of existing VLMs. During inference, when provided with user-specified region proposals, SpatialRGPT can accurately perceive their relative directions and distances. Additionally, we propose SpatialRGBT-Bench, a benchmark with ground-truth 3D annotations encompassing indoor, outdoor, and simulated environments, for evaluating 3D spatial cognition in VLMs. Our results demonstrate that SpatialRGPT significantly enhances performance in spatial reasoning tasks, both with and without local region prompts. The model also exhibits strong generalization capabilities, effectively reasoning about complex spatial relations and functioning as a region-aware dense reward annotator for robotic tasks. Code, dataset, and benchmark will be released at https://www.anjiecheng.me/SpatialRGPT.
@@ 1 Introduction
Recently, several works  [11, 17] has advanced VLMs’ spatial reasoning capabilities by introducing a comprehensive data generation pipeline that enables large-scale training with spatially-aware visual question answering (VQA) tasks. This approach is based on the hypothesis that the limited spatial reasoning capabilities of current VLMs are due to a lack of 3D/2D spatial knowledge in their training data. However, two critical challenges remain. First, effective spatial reasoning requires VLMs to accurately parse regional information, particularly the regions of object instances, whereas most existing VLMs are primarily designed to understand the global context of an image. When an image contains numerous instances, it becomes challenging to prompt the model to reason about the spatial relations between specific instances. This is because most VLMs function as global image parsers and do not support specifying regions for which users want to understand spatial relationships. Second, accurately perceiving spatial relations such as direction and distance cannot rely solely on RGB pixel data. Thus, the architecture needs to incorporate 3D inputs, such as depth information.
In this work, we propose SpatialRGPT, which leverages a data curation pipeline, along with a region and 3D-aware visual encoder architecture to significantly improve the spatial reasoning capability of VLMs.
@@
    We demonstrate downstream applications of SpatialRGPT. Leveraging SpatialRGPT’s region capabilities, we develop a region-aware dense reward annotator for robotics. Additionally, we show that SpatialRGPT can function as a stand-alone complex spatial reasoner.
@@ 3 Method
Our scene graph construction pipeline (Fig.1) begins with a filtering process to remove any unsuitable images (Appx.D.1). Using open-vocabulary models, we identify and ground candidate objects, followed by lifting them into 3D space using metric depth estimation and camera calibration. We then process the point clouds (Appx. D.3) to construct the final 3D scene graph.
@@
The 3D scene graph is a collection of tuples where the nodes represent specific 3D object instances, and the edges represent the spatial relationships between the nodes. Each node is defined by the object’s class, width, and height in metric scale. To create the node, we start by using the instance mask to deproject the object points from the depth map. Then, we perform canonicalization and denoising, and build 3D axis-aligned bounding boxes for each object. With the 3D bounding box, we can calculate the width and height of the objects in real-world units. The edges represent the spatial relationships between the nodes, including two types of relations: relative and metric. Relative relations contain left, right, above, below, behind, front, wide, thin, tall, short, big, and small. Metric relations include direction, direct distance, horizontal distance, and vertical distance between the two objects. We then traverse all the object nodes and use the point cloud centroids and bounding boxes to calculate their spatial relationships.
@@ 3.2 Learning Spatial-aware VLMs from 3D Scene Graph
In this section, we discuss converting the constructed 3D scene graph into textual representations for VLM training. One simple approach is through template-based methods via predefined handcrafted instructions. However, this approach limits the diversity of instructions and hinder the model’s reasoning capabilities. Thus, we employ additional complex QAs to enhance the model’s reasoning ability. Our results in Fig. 4 show that blending these two types of data can lead to a generalized and complex spatial reasoning model.
@@
These QAs serve as the foundation for learning basic spatial knowledge. We extract information about node attributes such as width and height, as well as relative and metric relations from the edge attributes. We create both qualitative and quantitative templates to generate questions and answers for each type of attribute, using entities in the form of Region [X]. This approach results in examples shown in the first row of Fig. 2. We provide detailed templates for each attribute in Appx. D.4.
@@
We employ Llama3-70B to generate complex spatial reasoning questions to enhance the model’s spatial reasoning capabilities. One approach is to input the scene graph directly into the LLMs. However, LLMs struggle to utilize 3D coordinate information effectively [10], so we opt for an alternative approach. We first construct spatial descriptions in a language format. Similar to the template-based approach, we extract attributes from the scene graph and then construct template-based spatial descriptions based on these attributes. We combine the spatial descriptions and the region tags as inputs to the LLM. The LLM is then tasked with creating a complex reasoning question and answer that is based on the description and matches the context. Examples of LLM-generated QAs are shown in the second row of Fig. 2. Our LLM prompts for generating QAs are provided in Appx. D.5.
We use our automated annotation pipeline to annotate images from the OpenImages [48] dataset, which covers a wide range of subjects and is of high resolution. The resulting Open Spatial Dataset (OSD) contains 1M unique images and 5M open-vocabulary regions, each associated with a bounding box and segmentation mask. Furthermore, the dataset includes 8M template-based QAs and 700K LLM-based QAs. We will release our OSD dataset along with the implementation of our automated dataset annotation pipeline upon paper publication.
@@ 3.3 VLM Architecture
An overview of SpatialRGPT’s VLM architecture is shown in Fig. 3. SpatialRGPT consists of a visual encoder (Appx. E.1) to encode vision features, a region-feature extractor [18] to obtain region-level embeddings (Appx. E.2), linear connectors (Appx. E.3) to project multi-modal embeddings into the word embedding space, and a large language model using LLaMA2-7B for language processing. In this section, we will explain why and how we incorporate depth information into SpatialRGPT, as well as how SpatialRGPT handles tokenizations.
@@ 3.4 Training and Inference Paradigm
SpatialRGPT training includes three stages [49]: (i) Connector Feature Alignment, (ii) Visual Language Pre-training, and (iii) Visual Instruction-tuning. During the first stage, CC3M image-caption pairs are used to pretrain the RGB connector as [28, 50, 51]. In the second stage, the visual language corpus from MMC4 [52] and COYO [53] is used to pretrain the LLM and the RGB connector. The RGB connector and LLM parameters are then frozen, with only the depth connector trainable and pre-trained on our OSD dataset. Finally, at stage three, we fine-tune the pre-trained model on visual language instruction-following datasets, using a combination of the instruction tuning dataset from [28], region-level instruction tuning data [18], and our OSD dataset. Detailed data blend of the visual instruction data is in Appx. F.1. For training region-level data and our OSD, we randomly sample from different modalities (e.g., box, mask) for each sample to ensure the model is versatile to the input modality. At inference time, SpatialRGPT can take both boxes or masks as input. For the results shown in the main paper, if the segmentation is available, we use the mask; if not, we use the box provided and apply SAM to segment the corresponding mask.
@@ 4 Experiments
RegionVILA [18] 30.83 47.61 35.84 44.64 35.57 49.09 40.48
SpatialRGPT 99.17 99.04 79.24 89.28 83.65 87.27 89.80
SpatialRGPT-Depth 99.17 99.04 80.19 91.96 87.50 91.81 91.78
@@
LLaVA-v1.6-34B [55]+SoM [56] 12.8 / 1.15 20.4 / 1.79 11.3 / 0.95 9.02 / 0.91 7.52 / 3.11 12.8 / 33.3°
KOSMOS-2 [8] 4.05 / >10 4.91 / >10 1.89 / 2.26 3.01 / 5.42 1.50 / 3.82 1.86 / 104°
RegionVILA [18] 22.3 / 1.30 24.6 / 3.26 17.9 / >10 36.8 / >10 49.6 / 1.61 35.5 / 79.8°
SpatialRGPT 35.1 / 0.35 59.0 / 0.27 53.8 / 0.27 51.9 / 0.31 54.9 / 0.63 95.3 / 17.1°
SpatialRGPT-Depth 41.2 / 0.33 65.6 / 0.25 51.9 / 0.27 49.6 / 0.31 57.9 / 0.61 95.3 / 15.4°
@@
Currently, there are no visual-language benchmarks that specifically focus on VLM’s ability to understand 3D spatial concepts like metric distance or size differences between objects. Recently, SpatialVLM created a spatial reasoning VQA benchmark using human labelers to annotate spatial information on 2D images, but this benchmark is not publicly available. To address this gap, we develop SpatialRGPT-Bench, a spatial reasoning VQA benchmark using data from both urban (nuScenes [57], KITTI [58]) and indoor (SUNRGBD [59], ARKitScenes [60]) environments, as well as simulated scenes (Hypersim [61]). These datasets cover various potential applications and include diverse object types, enhancing our benchmark’s thoroughness. We use preprocessed 3D cuboids for each object from Omni3D [62], all positioned within a unified 3D camera coordinate system and categorized by object classes. With these 3D cuboid annotations, we developed a conversational benchmark using our data generation pipeline. This benchmark comprises 657 qualitative and 749 quantitative VQA pairs, covering 88 distinct classes. All the samples come from the validation or test splits of the original datasets and are unseen by SpatialRGPT during the training phase. Please see Appx. B for statistics and examples of SpatialRGPT-Bench.
@@
[…]
VILA-7B [49] 79.9 62.3 57.8 68.2 64.4 85.5 34.9 68.9
RegionVILA-7B [18] 80.4 61.9 55.9 69.9 63.6 84.3 33.3 69.0
SpatialRGPT 80.3 62.5 58.2 67.8 63.4 85.3 31.5 68.6
SpatialRGPT-Depth 80.3 62.7 52.7 66.6 62.8 85.4 33.9 67.9
Table 2: SpatialRGPT results on Vision-language Benchmarks.
@@
Region-aware VLMs. This category explores models with region-level capabilities similar to our method. The models do not receive any language captions or object class information related to the region of interest; they rely solely on their visual processing capabilities. We equip GPT-4V [54] and LLaVA-v1.6-34B with Set of Marks (SoM) [56] to enable region-referring capabilities. Additionally, we include KOSMOS-2 [23], a VLM capable of taking bounding box inputs to reference objects, and RegionVILA (RegionGPT [18] with VILA [49] pre-training). RegionVILA also serves as an ablation baseline to our method; it shares the same model architecture as our RGB variant but is trained without our specialized spatial VQA dataset.
@@
4.2 Standard Visual-language Benchmarks
In this section, we evaluate whether integrating spatial VQA data and depth information affects performance on other VQA tasks. We compared our models with VILA-7B, which is trained on general VQA datasets, and RegionVILA [18], which includes training on both general and region-level datasets.
As shown in Table 2, our variants performed similarly to the baselines and slightly better on the VQA-v2 and GQA datasets. These results align with findings from [17], indicating that VLMs generally underperform on spatial reasoning tasks but can improve with specific spatial VQA training without compromising general VQA performance. This highlights the value of our data curation pipeline.
@@ 4.3 Real-world Applications
In this application, we aim to explore whether SpatialRGPT can function as a complex spatial reasoner on its own. Unlike the system mentioned in [17], which uses GPT-4 to handle reasoning tasks and employs VLM solely for answering basic spatial queries, SpatialRGPT directly integrates these capabilities. We provide examples in Fig 4, where we compare SpatialRGPT’s responses to those from GPT-4V using real-world samples. Our model demonstrates the ability to address complex spatial questions based on its own spatial knowledge. This suggests that SpatialRGPT has developed a robust representation of spatial learning and that this knowledge has effectively generalized to enhance its intrinsic language reasoning abilities.
@@
Recently, [17] has shown that VLMs can function as dense reward annotators for robotics tasks by specifying tasks in natural language and having the model annotate rewards for each frame in a trajectory. However, this approach can be constrained by the language’s ambiguity, especially when multiple identical objects are present or when targeting a small, specific region in a scene, which can be difficult to describe precisely with language alone. Given that SpatialRGPT is equipped with region-aware capabilities, we can directly specify the regions of interest. To study this application, we conducted a real robot experiment. Specifically, we defined two regions using bounding boxes (one for the fingertip and one for a green cube) and tasked SpatialRGPT to annotate rewards using the distance between the two regions. The results, shown in Fig. 1, indicate that the estimated distance between the fingertip and its target cube decreased monotonically as the fingertip moved towards its goal. Also, our depth variant performs slightly better than the RGB variant. This demonstrates SpatialRGPT ’s effectiveness as a region-aware dense reward annotator, offering a more precise and efficient alternative to language-only approaches.

### v3

@@ Abstract
Vision Language Models (VLMs) have demonstrated remarkable performance in 2D vision and language tasks. However, their ability to reason about spatial arrangements remains limited. In this work, we introduce Spatial Region GPT (SpatialRGPT) to enhance VLMs’ spatial perception and reasoning capabilities. SpatialRGPT advances VLMs’ spatial understanding through two key innovations: (i) a data curation pipeline that enables effective learning of regional representation from 3D scene graphs, and (ii) a flexible “plugin” module for integrating depth information into the visual encoder of existing VLMs. During inference, when provided with user-specified region proposals, SpatialRGPT can accurately perceive their relative directions and distances. Additionally, we propose SpatialRGBT-Bench, a benchmark with ground-truth 3D annotations encompassing indoor, outdoor, and simulated environments, for evaluating 3D spatial cognition in VLMs. Our results demonstrate that SpatialRGPT significantly enhances performance in spatial reasoning tasks, both with and without local region prompts. The model also exhibits strong generalization capabilities, effectively reasoning about complex spatial relations and functioning as a region-aware dense reward annotator for robotic tasks. Code, dataset, and benchmark are released at https://www.anjiecheng.me/SpatialRGPT.
@@ 1 Introduction
Recently, several works  [11, 17, 18] has advanced VLMs’ spatial reasoning capabilities by introducing a comprehensive data generation pipeline that enables large-scale training with spatially-aware visual question answering (VQA) tasks. This approach is based on the hypothesis that the limited spatial reasoning capabilities of current VLMs are due to a lack of 3D/2D spatial knowledge in their training data. However, two critical challenges remain. First, effective spatial reasoning requires VLMs to accurately parse regional information, particularly the regions of object instances, whereas most existing VLMs are primarily designed to understand the global context of an image. When an image contains numerous instances, it becomes challenging to prompt the model to reason about the spatial relations between specific instances. This is because most VLMs function as global image parsers and do not support specifying regions for which users want to understand spatial relationships. Second, accurately perceiving spatial relations such as direction and distance cannot rely solely on RGB pixel data. Thus, the architecture needs to incorporate 3D inputs, such as depth information.
In this work, we propose SpatialRGPT, leveraging a data curation pipeline, along with a region and 3D-aware visual encoder architecture to improve the spatial reasoning capability of VLMs.
@@
    We demonstrate downstream applications of SpatialRGPT. Leveraging SpatialRGPT’s region capabilities, we develop a region-aware dense reward annotator for robotics. Additionally, we show that SpatialRGPT can function as a stand-alone complex spatial reasoner, as well as its capacity to perform multi-hop reasoning.
@@ 3 Method
Our scene graph construction pipeline (Figure1) begins with a filtering process to remove any unsuitable images (Appx.F.1). Using open-vocabulary models, we identify and ground candidate objects, followed by lifting them into 3D space using metric depth estimation and camera calibration. We then process the point clouds (Appx. F.3) to construct the final 3D scene graph.
@@
The 3D scene graph is a collection of tuples where the nodes represent specific 3D object instances, and the edges represent the spatial relationships between the nodes. Each node is defined by the object’s class, width, and height in metric scale. To create the node, we start by using the instance mask to deproject the object points from the depth map. Then, we perform canonicalization and denoising, and build 3D axis-aligned bounding boxes for each object. With the 3D bounding box, we calculate the width and height of the objects in real-world units. The edges represent the spatial relationships between the nodes within two types of relations: relative and metric. Relative relations contain left, right, above, below, behind, front, wide, thin, tall, short, big, and small. Metric relations include direction, direct distance, horizontal distance, and vertical distance between the two objects. We then traverse all the object nodes and use the point cloud centroids and bounding boxes to calculate their spatial relationships.
@@ 3.2 Learning Spatial-aware VLMs from 3D Scene Graph
In this section, we discuss converting the constructed 3D scene graph into textual representations for VLM training. One simple approach is through template-based methods via predefined handcrafted instructions. However, this approach limits the diversity of instructions and hinder the model’s reasoning capabilities. Thus, we employ additional complex QAs to enhance the model’s reasoning ability. Our results in Figure 4 show that blending these two types of data can lead to a generalized and complex spatial reasoning model.
@@
These QAs serve as the foundation for learning basic spatial knowledge. We extract information about node attributes such as width and height, as well as relative and metric relations from the edge attributes. We create both qualitative and quantitative templates to generate questions and answers for each type of attribute, using entities in the form of Region [X]. This approach results in examples shown in the first row of Figure 2. We provide detailed templates for each attribute in Appx. F.4.
@@
We employ Llama3-70B to generate complex spatial reasoning questions to enhance the model’s spatial reasoning capabilities. One approach is to input the scene graph directly into the LLMs. However, LLMs struggle to utilize 3D coordinate information effectively [10], so we opt for an alternative approach. We first construct spatial descriptions in a language format. Similar to the template-based approach, we extract attributes from the scene graph and then construct template-based spatial descriptions based on these attributes. We combine the spatial descriptions and the region tags as inputs to the LLM. The LLM is then tasked with creating a complex reasoning question and answer that is based on the description and matches the context. Examples of LLM-generated QAs are shown in the second row of Figure 2. Our LLM prompts for generating QAs are provided in Appx. F.5.
We use our automated annotation pipeline to annotate images from the OpenImages [49] dataset, which covers a wide range of subjects and is of high resolution. The resulting Open Spatial Dataset (OSD) contains 1M unique images and 5M open-vocabulary regions, each associated with a bounding box and segmentation mask. Furthermore, the dataset includes 8M template-based QAs and 700K LLM-based QAs.
@@ 3.3 VLM Architecture
An overview of SpatialRGPT’s VLM architecture is shown in Figure 3. SpatialRGPT consists of a visual encoder (Appx. G.1) to encode vision features, a region-feature extractor [19] to obtain region-level embeddings (Appx. G.2), linear connectors (Appx. G.3) to project multi-modal embeddings into the word embedding space, and a large language model using LLaMA2-7B for language processing. In this section, we will explain why and how we incorporate depth information into SpatialRGPT, as well as how SpatialRGPT handles tokenizations.
@@ 3.4 Training and Inference Paradigm
SpatialRGPT training includes three stages [50]: (i) Connector Feature Alignment, (ii) Visual Language Pre-training, and (iii) Visual Instruction-tuning. During the first stage, CC3M image-caption pairs are used to pretrain the RGB connector as [29, 51, 52]. In the second stage, the visual language corpus from MMC4 [53] and COYO [54] is used to pretrain the LLM and the RGB connector. The RGB connector and LLM parameters are then frozen, with only the depth connector trainable and pre-trained on our OSD dataset. Finally, at stage three, we fine-tune the pre-trained model on visual language instruction-following datasets, using a combination of the instruction tuning dataset from [29], region-level instruction tuning data [19], and our OSD dataset. Detailed data blend of the visual instruction data is in Appx. H.1. For training region-level data and our OSD, we randomly sample from different modalities (e.g., box, mask) for each sample to ensure the model is versatile to the input modality. At inference time, SpatialRGPT can take both boxes or masks as input. For the results shown in the main paper, if the segmentation is available, we use the mask; if not, we use the box provided and apply SAM to segment the corresponding mask.
@@ 4 Experiments
RegionVILA-7B [19] 30.83 47.61 35.84 44.64 35.57 49.09 40.48
[…]
SpatialRGPT-7B 99.17 99.04 80.19 91.96 87.50 91.81 91.78
SpatialRGPT-VILA-1.5-3B 99.17 100.0 81.13 88.39 85.57 93.63 91.47
SpatialRGPT-VILA-1.5-8B 99.17 100.0 84.90 89.28 91.34 90.90 92.69
@@
LLaVA-v1.6-34B [56]+SoM [57] 12.8 / 1.15 20.4 / 1.79 11.3 / 0.95 9.0 / 0.91 7.5 / 3.11 12.8 / 33.3°
KOSMOS-2 [8] 4.1 / >10 4.91 / >10 1.9 / 2.26 3.0 / 5.42 1.5 / 3.82 1.9 / 104°
RegionVILA-7B [19] 22.3 / 1.30 24.6 / 3.26 17.9 / >10 36.8 / >10 49.6 / 1.61 35.5 / 79.8°
[…]
SpatialRGPT-7B 41.2 / 0.33 65.6 / 0.25 51.9 / 0.27 49.6 / 0.31 57.9 / 0.61 95.3 / 15.4°
SpatialRGPT-VILA-1.5-3B 44.6 / 0.30 63.1 / 0.22 50.9 / 0.28 42.9 / 0.33 63.2 / 0.60 93.5 / 10.4°
SpatialRGPT-VILA-1.5-8B 45.9 / 0.31 68.0 / 0.22 56.6 / 0.28 48.9 / 0.28 61.7 / 0.41 95.3 / 9.7°
@@
Currently, there are no visual-language benchmarks that specifically focus on VLM’s ability to understand 3D spatial concepts like metric distance or size differences between objects. Recently, SpatialVLM created a spatial reasoning VQA benchmark using human labelers to annotate spatial information on 2D images, but this benchmark is not publicly available. To address this gap, we develop SpatialRGPT-Bench, a spatial reasoning VQA benchmark using data from both urban (nuScenes [58], KITTI [59]) and indoor (SUNRGBD [60], ARKitScenes [61]) environments, as well as simulated scenes (Hypersim [62]). These datasets cover various potential applications and include diverse object types, enhancing our benchmark’s thoroughness. We use preprocessed 3D cuboids for each object from Omni3D [63], all positioned within a unified 3D camera coordinate system and categorized by object classes. With these 3D cuboid annotations, we developed a conversational benchmark using our data generation pipeline. This benchmark comprises 657 qualitative and 749 quantitative VQA pairs, covering 88 distinct classes. All the samples come from the validation or test splits of the original datasets and are unseen by SpatialRGPT during the training phase. Please see Appx. E for statistics and examples of SpatialRGPT-Bench.
@@
VQAv2 GQA SQAI VQAT POPE MME MMB
VILA-1.5-3B 80.4 61.5 69.0 60.4 85.9 1442 63.4
SpatialRGPT-VILA-1.5-3B 81.1 62.3 71.0 61.7 85.5 1424 65.6
MMB-CN SEED SEEDI MMMUV MMMUT LLaVAB MMVet
VILA-1.5-3B 52.7 60.9 67.9 33.3 30.8 75.9 35.4
SpatialRGPT-VILA-1.5-3B 53.6 61.8 69.0 33.0 31.3 71.5 38.2
Table 2: Comparison of SpatialRGPT and base model performance on general VLM benchmarks.
@@
[…]
4.2 Public Vision-language Benchmarks
General Benchmarks.
In this section, we evaluate whether integrating spatial VQA data and depth information affects performance on other VQA tasks. We compared our models with VILA-1.5-3B, which is trained on general VQA datasets. As shown in Table 2, our variants performed similarly to the baselines and slightly better on the VQA-v2 and MMVet datasets. These results align with findings from [17], indicating that VLMs generally underperform on spatial reasoning tasks but can improve with specific spatial VQA training without compromising general VQA performance.
Region & Spatial Benchmarks.
We follow the evaluation protocol from RegionGPT [19] and report object classification results using ground-truth boxes on the COCO-2017 validation set. As shown in Table 4, SpatialRGPT outperforms the baselines, demonstrating its strong region cognition capabilities. We further evaluate SpatialRGPT on BLINK [9]’s Relative Depth Benchmark. This benchmark is particularly challenging as it assesses point-level depths, while both the point-level region input and point-level questions were not specifically included in the training of SpatialRGPT. We use bounding boxes to mark the target points and evaluate the test set online with the EvalAI server. As shown in Table 4, SpatialRGPT significantly outperforms the state-of-the-art, achieving over 20% accuracy gain compared to GPT-4V-Turbo. Our model demonstrated strong performance, highlighting its ability to generalize to new tasks without explicit training.
Model mAP ($\uparrow$) Acc. (%)
CLIP [64] 58.9 -
RegionCLIP [65] 58.3 -
LLaVA-7B [29] - 40.0
Shikra-7B [25] - 53.9
GPT4RoI-7B [31] - 64.0
PVIT-7B [66] - 64.5
ASM-7B [32] 69.3 -
RegionGPT-7B [19] 70.0 80.6
SpatialRGPT-7B 69.7 79.9
SpatialRGPT-VILA-1.5-3B 72.5 82.5
SpatialRGPT-VILA-1.5-8B 72.9 82.9
Table 3: Region-level classification results. We follow the evaluation in RegionCLIP [65] and RegionGPT [19], report the results of object classification with ground-truth box on COCO-2017 validation set.
     Model      Acc. (%)
     Qwen-VL-Max [67]      58.9
     Gemini Pro [68]      50.0
     Claude 3 OPUS [69]      57.3
     GPT-4V-preview [55]      58.9
     GPT-4V-Turbo [55]      66.9
     GPT-4o [55]      64.5
     InstructBLIP-13B [51]      50.0
     Yi-VL-34B [70]      53.2
     LLaVA-v1.5-13B-xtuner [71]      54.0
     LLaVA-v1.6-34B [56]      64.5
     MiniGPT-4-v2-7B [26]      49.2
     InstructBLIP-7B [51]      50.8
     LLaVA-v1.5-7B-xtuner [71]      50.8
     CogVLM-7B [27]      50.8
     LLaVA-v1.5-7B [72]      51.6
     LLaVA-internLM2-7B [73]      52.4
     SpatialRGPT-7B      82.3
     SpatialRGPT-VILA-1.5-8B      87.9
Table 4: BLINKRelativeDepth results.
Figure 5: Examples of SpatialRGPT performing multi-hop reasoning.
@@ 4.3 Real-world Applications
In this application, we aim to explore whether SpatialRGPT can function as a complex spatial reasoner on its own. Unlike the system mentioned in [17], which uses GPT-4 to handle reasoning tasks and employs VLM solely for answering basic spatial queries, SpatialRGPT directly integrates these capabilities. We provide examples in Figure 4, where we compare SpatialRGPT’s responses to those from GPT-4V using real-world samples. Our model demonstrates the ability to address complex spatial questions based on its own spatial knowledge. This suggests that SpatialRGPT has developed a robust representation of spatial learning and that this knowledge has effectively generalized to enhance its intrinsic language reasoning abilities.
Multi-hop Reasoning.
In Figure 5, we show examples of SpatialRGPT handling multi-hop reasoning. In the upper left sample, the model first identifies what’s to the right of Region [0] (a single apple), finds the basket there, determines what’s inside the basket, and then provides spatial details about the object inside. Even though our training data doesn’t specifically include such multi-hop tasks, SpatialRGPT can still manage them effectively. This indicates that the model has developed a strong understanding of spatial relationships.
@@
Recently, [17] has shown that VLMs can function as dense reward annotators for robotics tasks by specifying tasks in natural language and having the model annotate rewards for each frame in a trajectory. However, this approach can be constrained by the language’s ambiguity, especially when multiple identical objects are present or when targeting a small, specific region in a scene, which can be difficult to describe precisely with language alone. Given that SpatialRGPT is equipped with region-aware capabilities, we can directly specify the regions of interest. To study this application, we conducted a real robot experiment. Specifically, we defined two regions using bounding boxes (one for the fingertip and one for a green cube) and tasked SpatialRGPT to annotate rewards using the distance between the two regions. The results, shown in Figure 6, indicate that the estimated distance between the fingertip and its target cube decreased monotonically as the fingertip moved towards its goal. Also, our depth variant performs slightly better than the RGB variant. This demonstrates SpatialRGPT ’s effectiveness as a region-aware dense reward annotator, offering a more precise and efficient alternative to language-only approaches.